OpenStudy (anonymous):

What is the solution to the rational equation 2x/3 + 4/3 = 6? a. x = 4 b. x = 5 c. x = 7 d. x = 11 I think its c... am i right?

OpenStudy (anonymous):

yup....

jhonyy9 (jhonyy9):

yes you are right

jhonyy9 (jhonyy9):

but can you writing here these calcules ?

OpenStudy (anonymous):

whenever you want to check your answer substitute your solution back into the original equation, eg if you think its c but you aren't sure: 2(7)/3 + 4/3 = 6 14/3 + 4/3 = 18/3 = 6 so you are indeed correct

OpenStudy (anonymous):

thanks!